OpenStudy (anonymous):

Write the sum 384 +242 in expanded form.

OpenStudy (anonymous):

example of expanded form 123,456,789 = 100,000,000 + 20,000,000+3,000,000+400,000+50,000+6,000+700+80+9

OpenStudy (anonymous):

how did u get that

OpenStudy (anonymous):

thats nothing like my options

OpenStudy (anonymous):

600+20+6

OpenStudy (anonymous):

im kinda not sure

OpenStudy (akashdeepdeb):

384 +242 = 626 Hundreds Tens Ones 6 2 6 So the expanded form is 6*100 + 2*10 + 6*1 600 + 20 + 6 Understood? :)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

little lost

OpenStudy (anonymous):

these are my choices for my answers

OpenStudy (anonymous):

Do you have options so we can get which type of answer you want?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Write the sum in expanded form. A) B) C) D) E) 4.

OpenStudy (anonymous):

hang on trying to copy over to you

OpenStudy (anonymous):

that is the example i need it written

OpenStudy (anonymous):

okie tell me what is ans of 384 +242

OpenStudy (akashdeepdeb):

First of all it should not be a 12 and should be a 2 Second of all, I wrote the exact same thing.

OpenStudy (anonymous):

its 626 so what you do is write it as 600 +20+6 now 600 = 6*100 = 6*10^2 similarly 20=2*10= 2*10^1 and last of all 6=6*1=6*10^0 so there you go

OpenStudy (anonymous):

ok I was wondering how to get that answer

OpenStudy (anonymous):

kind of wanted sone to explain it

OpenStudy (anonymous):

tell me what you didnt get in the answer?

OpenStudy (anonymous):

i was just trying to figure it out but i got it now. thank you

OpenStudy (anonymous):

?